# Trivial Isomorphisms between Reduced Products

### v3

@@ 4. Automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$
is a C-measurable lifting of $\Phi$ on ${\mathcal{J}}_{\textrm{cont}}$. By [8, Lemma 3.3.4], $\Phi$ decomposes as $\Phi_{1}\oplus\Phi_{2}$ where $\Phi_{1}$ has a continuous lifting and the kernel of $\Phi_{2}$ is nonmeager. By [8, Theorem 1.6.1], $\Phi_{1}$ has a completely additive lifting, as required. ∎
The conclusion of Theorem 4.7 is incompatible with CH. An obvious route towards strengthening it is by requiring $\Phi$ to have a continuous lifting; this strengthening was even named Strong Extension Principle and conjectured to follow from forcing axioms in [8, §4] (the Weak Extension Principle is a generalization of the conclusion of Theorem 4.7—or rather, its dual restatement in terms of Boolean algebras, see [12, §1.3.1]—to Čech–Stone remainders of locally compact Polish spaces other than ${\mathbb{N}}$). This conjecture is, however, false. By [5], there is an endomorphism of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ which does not have a lifting which is an endomorphism of ${\mathcal{P}}({\mathbb{N}})$.
@@ 5. Proof of Theorem 7
We will prove that for every tree-like almost disjoint family ${\mathcal{A}}$ all but countably many of its elements belong to ${\mathcal{J}}_{\rm{prod}}$ (see (5.1)). Using the well-known characterization of nonmeagerness of ideals on ${\mathbb{N}}$ due to Jalali-Naini and Talagrand ([8, Theorem 3.10.1]) again, this will imply that the ideal ${\mathcal{J}}_{\rm{prod}}$ is nonmeager.
@@ 5.2. Uniformization
Consider the Polish space ${\mathcal{Y}}_{n}=({\mathcal{N}}_{n})^{{\mathcal{M}}_{n}}$. Then $\{(g^{A},A)\mid A\in{\mathcal{J}}_{\rm{prod}}\}$ is a family of partial selectors as defined and studied in §3.1. The following shows that the functions $g^{A}$, for $A\in{\mathcal{J}}_{\rm{prod}}$, moreover form a coherent family in the sense of [8, §2].
@@
Proposition 5.2 implies that ${\mathcal{J}}_{\rm{prod}}$ is nonmeager (recall that this conclusion is included in the assumptions of Proposition 5.6), and (again by Talagrand and Jalali-Naini, [8, Theorem 3.10.1]) it intersects every perfect tree-like almost disjoint family. Fix $B\in{\mathcal{A}}\cap{\mathcal{J}}_{\rm{prod}}$. For $x\in Z$ let $n(x)=\max(\operatorname{Diff}(g^{f(x)},g^{B}))$ (with $\max\emptyset=0$). Since $Z$ is uncountable, there is $n$ such that $Z^{\prime}=\{x\in Z\mid n(x)<n\}$ is uncountable. Choose distinct $y,z$ in $Z^{\prime}$ satisfying $\Delta(y,z)\geq n$ and fix $j$ in $(\operatorname{Diff}(g^{f(y)},g^{f(z)})\cap B)\setminus n$. Then $g^{B}_{j}=g^{f(y)}_{j}\neq g^{f(z)}_{j}=g^{B}_{j}$; contradiction. ∎
@@
[8, Theorem 3.10.1(a)] implies that there exists $\bar{k}$ such that every finite subset of ${\mathbb{N}}\setminus\bar{k}$ is in $\hat{\mathcal{G}}_{\bar{n}}$. Let
@@
Assume otherwise and fix $(a,b)\in{\mathcal{Y}}$ such that $S=\{n\mid b(n)\neq\Phi_{*}(a)(n)\}$ is infinite. Since $\hat{\mathcal{G}}_{\bar{n}}$ is nonmeager and hereditary, there is an infinite $X\subseteq S$ in $\hat{\mathcal{G}}_{\bar{n}}$ ([8, Theorem 3.10.1 (a)] gives a considerably stronger statement). Let $Y\in{\mathcal{G}}_{\bar{n}}$ be such that $X\subseteq Y$ and note that $\Psi_{Y}(a)=^{*}\Phi_{*}(a)\restriction Y$. Let $l\geq\bar{k}$ be such that $g^{Y}_{i}(a(i))=\Phi_{*}(a)(i)$ for all $i\geq l$ and find $m$ such that $X_{0}=X\cap[l,m)$ has at least $\bar{n}+1$ elements. By the definition of ${\mathcal{Y}}$, there is $C\in{\mathcal{G}}_{\bar{n}}$ such that $g^{C}_{i}(a(i))=b(i)$ for all $i\in X_{0}$. This implies $\Delta(g^{C},g^{Y})\supseteq X_{0}$. Since $|X_{0}|\geq\bar{n}+1$, this contradicts the homogeneity property of ${\mathcal{G}}_{\bar{n}}$ as in Claim 5.8. ∎
@@
are equal to $\Phi_{*}(a)$ modulo finite, in particular ${\mathcal{Y}}_{a}$ is countable. By the Luzin–Novikov Theorem ([26, Theorem 18.10]), there is a Borel-measurable uniformization $\Theta\colon\prod_{n}{\mathcal{M}}_{n}\to\prod_{n}{\mathcal{N}}_{n}$. Claim 5.9 implies that $\Theta$ is a lifting of $\Phi$. ∎
@@ 5.3. Automorphisms with Baire-measurable liftings
Consider first the case when all ${\mathcal{M}}_{n}$ are finite. In this case we will not need $\operatorname{{\mathsf{OCA}_{T}}}$. Since $\Theta$ is Baire-measurable, it is continuous on a dense $G_{\delta}$ set $G$. By the standard representation of dense $G_{\delta}$ sets in $\prod_{n}{\mathcal{M}}_{n}$ (see e.g., [12, Theorem 9.9.1]) there are a partition ${\mathbb{N}}=S_{0}\sqcup S_{1}$, $a_{0}\in\prod_{n\in S_{0}}{\mathcal{M}}_{n}$, and $a_{1}\in\prod_{n\in S_{1}}{\mathcal{M}}_{n}$ such that for $j\in\{0,1\}$ we have $\{b\in\prod_{n}{\mathcal{M}}_{n}\mid b\restriction S_{j}=a_{j}\}\subseteq G$. (This is precisely where the finiteness of ${\mathcal{M}}_{n}$ gets used, as this characterisation does not hold for infinite sets). If $x\in\prod_{n\in S_{0}}{\mathcal{M}}_{n}$ and $y\in\prod_{n\in S_{1}}{\mathcal{M}}_{n}$, let $x^{\frown}y$ be the concatenation of $x$ and $y$, that is, the unique $z\in\prod_{n}{\mathcal{M}}_{n}$ which equals $x$ on $S_{0}$ and $y$ on $S_{1}$.
@@
It is very likely that the version of Proposition 5.10 in which $\Theta$ is assumed to be Borel-measurable can be proven in $\operatorname{{\mathsf{ZFC}}}$, using the fact that $\operatorname{{\mathsf{OCA}_{T}}}$ for analytic sets is a theorem of $\operatorname{{\mathsf{ZFC}}}$ ([16]). A similar argument was used in [8, Theorem 1.6.1 and 1.6.2].
@@ 8. Concluding remarks
First, what other theories recognize coordinates in reduced products? Can recognizing coordinates be re-framed in terms of the existence of a copy of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ in the $\operatorname{eq}$ (see [23, p. 151]) of the reduced product at hand? Or, are theories that recognize coordinates characterized by the existence of local copies of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$, like the ones obtained for ramified sets and random graphs (§2.2 and §2.3), in reduced products of its models? Long story short, is there a natural dividing line for theories $T$ such that the reduced products of models of $T$ recognize coordinates (compare with the order property/stability as in [32])? An easy observation is that if $T$ recognizes coordinates and $T’$ ’ is a theory in an expanded language extending $T$, then $T’$ ’ recognizes coordinates as well.
@@ Appendix A From $\sigma$-Borel to continuous
Suppose that $X$ and $Y$ are Polish spaces. A function $\Theta\colon X\to Y$ is called C-measurable if it is measurable with respect to the $\sigma$-algebra generated by analytic sets (sometimes denoted $\sigma(\Sigma^{1}_{1})$). The following is a consequence of Jankov–von Neumann’s uniformization theorem (see [26, 18.A]):
@@
is, by Novikov’s theorem ([26, Theorem 29.3]), analytic. By our assumption, for every $a\in{\mathcal{T}}$ the set ${\mathcal{Z}}(a)=\{b\subseteq B\mid(a,b,\Phi_{*}(a))\in{\mathcal{X}}\}$ is comeager in ${\mathcal{P}}(B)$, in particular the section ${\mathcal{Y}}_{a}$ is nonempty for all $a\in{\mathcal{T}}$.
@@
Suppose that $a_{n}$, ${\mathcal{X}}_{n}$, $(U_{ni})_{i}$ and $[s_{n}]$ as required had been chosen. Then we can write $[s_{n}]\cap{\mathcal{P}}(C_{n})=[t_{n}]\cap{\mathcal{P}}(A_{n+1})\oplus[u_{n}]\cap{\mathcal{P}}(C_{n+1})$ for clopen sets $[t_{n}]\cap{\mathcal{P}}(A_{n+1})\subseteq{\mathcal{P}}(A_{n+1})$ and $[u_{n}]\cap{\mathcal{P}}(C_{n+1})\subseteq{\mathcal{P}}(C_{n+1})$. By the Kuratowski–Ulam theorem (see e.g., [26, §8.K]), the set

### v4

@@ 4. Automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$
is a C-measurable lifting of $\Phi$ on ${\mathcal{J}}_{\textrm{cont}}$. By [9, Lemma 3.3.4], $\Phi$ decomposes as $\Phi_{1}\oplus\Phi_{2}$ where $\Phi_{1}$ has a continuous lifting and the kernel of $\Phi_{2}$ is nonmeager. By [9, Theorem 1.6.1], $\Phi_{1}$ has a completely additive lifting, as required. ∎
The conclusion of Theorem 4.7 is incompatible with CH. An obvious route towards strengthening it is by requiring $\Phi$ to have a continuous lifting; this strengthening was even named Strong Extension Principle and conjectured to follow from forcing axioms in [9, §4] (the Weak Extension Principle is a generalization of the conclusion of Theorem 4.7—or rather, its dual restatement in terms of Boolean algebras, see [13, §1.3.1]—to Čech–Stone remainders of locally compact Polish spaces other than ${\mathbb{N}}$). This conjecture is, however, false. By [6], there is an endomorphism of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ which does not have a lifting which is an endomorphism of ${\mathcal{P}}({\mathbb{N}})$.
@@ 5. Proof of Theorem 7
We will prove that for every tree-like almost disjoint family ${\mathcal{A}}$ all but countably many of its elements belong to ${\mathcal{J}}_{\rm{prod}}$ (see (5.1)). Using the well-known characterization of nonmeagerness of ideals on ${\mathbb{N}}$ due to Jalali-Naini and Talagrand ([9, Theorem 3.10.1]) again, this will imply that the ideal ${\mathcal{J}}_{\rm{prod}}$ is nonmeager.
@@ 5.2. Uniformization
Consider the Polish space ${\mathcal{Y}}_{n}=({\mathcal{N}}_{n})^{{\mathcal{M}}_{n}}$. Then $\{(g^{A},A)\mid A\in{\mathcal{J}}_{\rm{prod}}\}$ is a family of partial selectors as defined and studied in §3.1. The following shows that the functions $g^{A}$, for $A\in{\mathcal{J}}_{\rm{prod}}$, moreover form a coherent family in the sense of [9, §2].
@@
Proposition 5.2 implies that ${\mathcal{J}}_{\rm{prod}}$ is nonmeager (recall that this conclusion is included in the assumptions of Proposition 5.6), and (again by Talagrand and Jalali-Naini, [9, Theorem 3.10.1]) it intersects every perfect tree-like almost disjoint family. Fix $B\in{\mathcal{A}}\cap{\mathcal{J}}_{\rm{prod}}$. For $x\in Z$ let $n(x)=\max(\operatorname{Diff}(g^{f(x)},g^{B}))$ (with $\max\emptyset=0$). Since $Z$ is uncountable, there is $n$ such that $Z^{\prime}=\{x\in Z\mid n(x)<n\}$ is uncountable. Choose distinct $y,z$ in $Z^{\prime}$ satisfying $\Delta(y,z)\geq n$ and fix $j$ in $(\operatorname{Diff}(g^{f(y)},g^{f(z)})\cap B)\setminus n$. Then $g^{B}_{j}=g^{f(y)}_{j}\neq g^{f(z)}_{j}=g^{B}_{j}$; contradiction. ∎
@@
[9, Theorem 3.10.1(a)] implies that there exists $\bar{k}$ such that every finite subset of ${\mathbb{N}}\setminus\bar{k}$ is in $\hat{\mathcal{G}}_{\bar{n}}$. Let
@@
Assume otherwise and fix $(a,b)\in{\mathcal{Y}}$ such that $S=\{n\mid b(n)\neq\Phi_{*}(a)(n)\}$ is infinite. Since $\hat{\mathcal{G}}_{\bar{n}}$ is nonmeager and hereditary, there is an infinite $X\subseteq S$ in $\hat{\mathcal{G}}_{\bar{n}}$ ([9, Theorem 3.10.1 (a)] gives a considerably stronger statement). Let $Y\in{\mathcal{G}}_{\bar{n}}$ be such that $X\subseteq Y$ and note that $\Psi_{Y}(a)=^{*}\Phi_{*}(a)\restriction Y$. Let $l\geq\bar{k}$ be such that $g^{Y}_{i}(a(i))=\Phi_{*}(a)(i)$ for all $i\geq l$ and find $m$ such that $X_{0}=X\cap[l,m)$ has at least $\bar{n}+1$ elements. By the definition of ${\mathcal{Y}}$, there is $C\in{\mathcal{G}}_{\bar{n}}$ such that $g^{C}_{i}(a(i))=b(i)$ for all $i\in X_{0}$. This implies $\Delta(g^{C},g^{Y})\supseteq X_{0}$. Since $|X_{0}|\geq\bar{n}+1$, this contradicts the homogeneity property of ${\mathcal{G}}_{\bar{n}}$ as in Claim 5.8. ∎
@@
are equal to $\Phi_{*}(a)$ modulo finite, in particular ${\mathcal{Y}}_{a}$ is countable. By the Luzin–Novikov Theorem ([27, Theorem 18.10]), there is a Borel-measurable uniformization $\Theta\colon\prod_{n}{\mathcal{M}}_{n}\to\prod_{n}{\mathcal{N}}_{n}$. Claim 5.9 implies that $\Theta$ is a lifting of $\Phi$. ∎
@@ 5.3. Automorphisms with Baire-measurable liftings
Consider first the case when all ${\mathcal{M}}_{n}$ are finite. In this case we will not need $\operatorname{{\mathsf{OCA}_{T}}}$. Since $\Theta$ is Baire-measurable, it is continuous on a dense $G_{\delta}$ set $G$. By the standard representation of dense $G_{\delta}$ sets in $\prod_{n}{\mathcal{M}}_{n}$ (see e.g., [13, Theorem 9.9.1]) there are a partition ${\mathbb{N}}=S_{0}\sqcup S_{1}$, $a_{0}\in\prod_{n\in S_{0}}{\mathcal{M}}_{n}$, and $a_{1}\in\prod_{n\in S_{1}}{\mathcal{M}}_{n}$ such that for $j\in\{0,1\}$ we have $\{b\in\prod_{n}{\mathcal{M}}_{n}\mid b\restriction S_{j}=a_{j}\}\subseteq G$. (This is precisely where the finiteness of ${\mathcal{M}}_{n}$ gets used, as this characterisation does not hold for infinite sets). If $x\in\prod_{n\in S_{0}}{\mathcal{M}}_{n}$ and $y\in\prod_{n\in S_{1}}{\mathcal{M}}_{n}$, let $x^{\frown}y$ be the concatenation of $x$ and $y$, that is, the unique $z\in\prod_{n}{\mathcal{M}}_{n}$ which equals $x$ on $S_{0}$ and $y$ on $S_{1}$.
@@
It is very likely that the version of Proposition 5.10 in which $\Theta$ is assumed to be Borel-measurable can be proven in $\operatorname{{\mathsf{ZFC}}}$, using the fact that $\operatorname{{\mathsf{OCA}_{T}}}$ for analytic sets is a theorem of $\operatorname{{\mathsf{ZFC}}}$ ([17]). A similar argument was used in [9, Theorem 1.6.1 and 1.6.2].
@@ 8. Concluding remarks
First, what other theories recognize coordinates in reduced products? Can recognizing coordinates be re-framed in terms of the existence of a copy of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ in the $\operatorname{eq}$ (see [24, p. 151]) of the reduced product at hand? Or, are theories that recognize coordinates characterized by the existence of local copies of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$, like the ones obtained for ramified sets and random graphs (§2.2 and §2.3), in reduced products of its models? Long story short, is there a natural dividing line for theories $T$ such that the reduced products of models of $T$ recognize coordinates (compare with the order property/stability as in [33])? An easy observation is that if $T$ recognizes coordinates and $T’$ ’ is a theory in an expanded language extending $T$, then $T’$ ’ recognizes coordinates as well.
@@ Appendix A From $\sigma$-Borel to continuous
Suppose that $X$ and $Y$ are Polish spaces. A function $\Theta\colon X\to Y$ is called C-measurable if it is measurable with respect to the $\sigma$-algebra generated by analytic sets (sometimes denoted $\sigma(\Sigma^{1}_{1})$). The following is a consequence of Jankov–von Neumann’s uniformization theorem (see [27, 18.A]):
@@
is, by Novikov’s theorem ([27, Theorem 29.3]), analytic. By our assumption, for every $a\in{\mathcal{T}}$ the set ${\mathcal{Z}}(a)=\{b\subseteq B\mid(a,b,\Phi_{*}(a))\in{\mathcal{X}}\}$ is comeager in ${\mathcal{P}}(B)$, in particular the section ${\mathcal{Y}}_{a}$ is nonempty for all $a\in{\mathcal{T}}$.
@@
Suppose that $a_{n}$, ${\mathcal{X}}_{n}$, $(U_{ni})_{i}$ and $[s_{n}]$ as required had been chosen. Then we can write $[s_{n}]\cap{\mathcal{P}}(C_{n})=[t_{n}]\cap{\mathcal{P}}(A_{n+1})\oplus[u_{n}]\cap{\mathcal{P}}(C_{n+1})$ for clopen sets $[t_{n}]\cap{\mathcal{P}}(A_{n+1})\subseteq{\mathcal{P}}(A_{n+1})$ and $[u_{n}]\cap{\mathcal{P}}(C_{n+1})\subseteq{\mathcal{P}}(C_{n+1})$. By the Kuratowski–Ulam theorem (see e.g., [27, §8.K]), the set